OpenStudy (anonymous):

please help me solve -5+b over 4=7

OpenStudy (anonymous):

\[\frac{-5 + b}{4} = 7\] \[\frac{4}{1} * \frac{-5 + b}{4} = 7 * \frac{4}{1}\] \[-5 + b = 28\] \[b = 33\]